# Generative Data Mining with Longtail-Guided Diffusion

### v1

@@ Appendix A Appendix
Table 4: Top-1 Accuracy on ImageNet-LT by dataset split (Many, Med, Few) and overall (Acc). The validation set is balanced between all 1000 classes but trainig is highly imbalanced: Few classes have $5-19$ examples, Medium classes have $20-99$ examples, Many classes have $100-1200$ examples.
A.1 LTG Improves Longtail Performance
In Table 4, we train SOTA ViT-based models (LiVT) from scratch on ImageNet-LT according to the longtail-compensation approach of (Xu et al., 2023). In summary, training includes MAE pretraining followed by 100 epochs of BCE loss with a logit adjustment to account for class imbalance. Initial LiVT training is on real data only and uses many data augmentations including RandAugment, Mixup, and CutMix. Our baseline predictive model, LiVT (Reproduced), nearly matches the generalization performance (top-1 accuracy) of what is reported in (Xu et al., 2023) (LiVT (Reproduced): $60.6$ vs LiVT (Reported): $60.9$). We then fine-tune LiVT on $24000$ additional synthetic data evenly distributed across all $1000$ classes for $100$ epochs with $41e-$ 4 learning rate. We use the same Stable Diffusion 1.4 baseline and sampling details as in the main paper (LiVT SD), and also compare Longtail Guidance with three types of model-based longtail signals (LiVT LTG {Energy, Entropy, Epistemic}). Overall generalization improvements on the balanced validation set are mild but significant ($+1.3$ top-1 accuracy). However, performance for longtail classes (Few) improves by four points for baseline diffusion ($+10\%$) but a marked $10$ points for Longtail Guidance ($+24\%$)!
A.2 Comparison to GIF and Dream-ID
@@
In the main paper, Table 1, we compare to GIF, the Guided Imagination Framework (Zhang et al., 2023b). We use the same predictive model architecture (ResNet50 trained from scratch), generate the same quantity of synthetic data (30×30$\times$30 × dataset size for Pets, 20×20$\times$20 × dataset size for Caltech101, Cars, and Flowers), the same diffusion model (Stable Diffusion v1.4), the same diffusion sampler (DDIM), and the same number of sampling iterates ($50$). We reproduce baseline models (not exposed to synthetic data or data augmentation conditions) to match the performance of GIF baselines (Original row in Table 1) to within $1\%$ of the generalization performance reported in GIF. We then fine-tune for $100$ epochs, generating synthetic data with Longtail Guidance according to the schedule in Table 5 (e.g. generate synthetic data in epoch 0, fine-tune until epoch 5, generate more synthetic data, fine-tune until epoch 10, …). We experimentally found that synthetic data iteratively generated throughout fine-tuning outperformed synthetic data generated all at once and speculate that this is due to the model-based longtail evolving throughout training.
@@ A.4 Dataset Details
A.5 Can We Continuously Mine for Additional High-Value Synthetic Data?
@@ A.7 Longtail Signals
A.8 Why Does Longtail Guidance Work without Training the Predictive Model on Intermediate Diffusion States?
A key finding and contribution of this work is that an existing predictive model $f_{\phi}$ does not need to be retrained on intermediate, noisy diffusion states to effectively guide diffusion model $\epsilon_{\theta}$ towards high-value synthetic training examples that are rare or hard from $f$’s perspective. This realization frees us from the dilemma of having to decide whether to waste predictive model capacity training on intermediate diffusion data it will never see in production or risking divergence from the production model by fine-tuning on intermediate noisy states. But why does this work?
In Figure  15, we visualize the intermediate, noisy diffusion states of two quantities: the decoded, predicted terminal state $D(P(\hat{x}_{t}^{0}))$ (top row, what LTG uses as guidance input to predictive model $f_{\phi}$) and the decoded data state, $D(x_{t})$ (bottom row, what classifier guidance would traditionally performs guidance if not in a latent space). We observe that the terminal state predictions much more readily resemble natural image data at a much earlier time in the diffusion process (within the first 10% of denoising steps) than do the data states. In fact, decoded data states have off-distribution noise artifacts up through the first 90% of the diffusion denoising steps.
@@
Results are plotted in Figure 15. Observe that the decoded, predicted terminal state $D(P(\hat{x}_{t}^{0}))$ are dramatically closer (in distribution) to real training data than are the naively decoded data states, $x_{t}$ for nearly all denoising iterations.
@@ A.9 Ablations
A.10 Additional Examples
In Figure 18, we show high-resolution examples of synthetic data generated by baseline diffusion and by Longtail Guidance. In order, classes are: jellyfish, flamingo, lion, monarch butterfly, ant, candle, jeep, forklift, and ambulance. Guidance is performed by a SOTA ViT-based ImageNet-LT classifier described in Supplement A.1.
Figure 18: Additional example of baseline diffusion (left) and Longtail Guidance (right). Baseline diffusion tends to generate canonical, well-posed views that help a predictive generalize, but only up to a point. In contrast, Longtail Guidance produces more extreme, occluded, or challenging views from the perspective of a predictive model, enabling significantly improved generalization performance.

### v2

@@ 4.2 LTG Improves Longtail Performance
4.3 LTG Produces Meaningful Longtail Data
Refined prompts are used to generate additional synthetic data $x_{\text{inspect}}$ by diffusion (without LTG). We train $f_{\phi}$ on synthetic data generated by Longtail Introspection and compare it to synthetic data generated by manual prompt tuning (40 unique prompts per class, as in (Zhang et al., 2023b); details in Supplement A.2). Table 4 shows that Longtail Introspection significantly outperforms manually prompt-tuned diffusion (defined in Supplement A.1), quantitatively supporting that data generated by LTG exhibit meaningful variation. Figure 8 shows example longtail keywords.
Figure 8: Longtail keywords from LTG-generated Flowers data.
Condition Accuracy
No Synthetic Data 74.1
Manual Prompt Tuning 78.8
Longtail Introspection 84.3
4.4 Discussion
5 Related Work
Diffusion has seen wide success, particularly in image generation, where it outperform GANs in image quality and diversity without suffering from unstable training or mode collapse (Dhariwal & Nichol, 2021). Recent work has seen progress in handling large data dimensions with latent spaces (Chen et al., 2025; Podell et al., 2023; Rombach et al., 2022) or hourglass networks (Crowson et al., 2024), improved sampling (Lu et al., 2022; Karras et al., 2022; Ho et al., 2020; Song et al., 2020a), additional data domains (Ran et al., 2024; Pronovost et al., 2023; Zhong et al., 2023), and personalization (Ruiz et al., 2023; Kumari et al., 2023). Much work has also been done on new forms of guidance (Wallace et al., 2023b; Yu et al., 2023a; Wallace et al., 2023a; Zhang et al., 2023a) beyond just classifier guidance and classifier-free guidance (Dhariwal & Nichol, 2021; Ho & Salimans, 2022). Universal Guidance (Bansal et al., 2023) and Diffusion Posterior Sampling (Dou & Song, 2024) are most relevant and are discussed in Section 3.
Model signals have been used for out-of-distribution or adversarial detection (Huang et al., 2021; Hsu et al., 2020; Cohen et al., 2020), particularly model uncertainty (Van Amersfoort et al., 2020) or feature density (Lee et al., 2018). Epistemic uncertainty has been developed in expensive Bayesian (Gal & Ghahramani, 2016) or model ensemble contexts (Liu et al., 2019a; Depeweg et al., 2018; Wilson & Izmailov, 2020; Lakshminarayanan et al., 2017), though, to our knowledge, not in a single, differentiable forward pass as we have done. Work on longtail robustness has primarily focused on addressing a-priori known class imbalance. Datasets include ImageNet-LT, Places-LT, (Liu et al., 2019b) and iNaturalist (Van Horn et al., 2018). Mitigations include pretraining (He et al., 2022; Bao et al., 2021), distillation (Xiang et al., 2020), reweighted loss (Xu et al., 2023; Ross & Dollár, 2017), selective real data mining (Jiang et al., 2022), or low-density data sampling (Um & Ye, 2024).
Conclusion
We develop model-based longtail signals that do not impact model weights or performance and are leveraged by Longtail Guidance, a synthetic data generation approach that explicitly conditions diffusion on an existing predictive model to generate examples that are rare or hard from that model’s perspective. We show that training on LTG-generated data provides strong, data-efficient generalization improvements across eight datasets. We further demonstrate that longtail synthetic data generations can be rendered into meaningful text descriptions and keywords that can aid future (real or synthetic) data collection priorities.
Predictive models are being deployed more than ever before. Increasingly, they will encounter longtail scenarios that human operators cannot easily predict in advance. Foundation models can be used to mitigate some risk, but we cannot shoehorn the entirety of Internet-scale knowledge into every deployed model due to capacity and compute constraints. By letting an existing predictive model speak for itself, we can expend offline compute to generatively mine high-value synthetic training data. Further reducing these synthetic examples to human- and machine-readable text suggests a future where we can move away from slow, reactive longtail mitigation towards fast, proactive longtail discovery.
Impact Statement
We develop model-based longtail signals, model-conditioned longtail synthetic data generation, and longtail interpretation techniques that have the potential to discover and mitigate production model failure modes before they occur. This has tremendous potential for improving the quality and safety of deployed ML applications. However, they may also have negative impacts if human operators rely too heavily on our approaches and not enough on traditional safeguards. We recommend our methods be used to supplement, not replace, standard deployment practices; most importantly, that regular and satisfactory evaluation on non-synthetic evaluation data be used as a requirement for model deployment.
Appendix A Appendix
A.1 Comparison to GIF and Dream-ID
@@
In the main paper, Table 1, we compare to GIF, the Guided Imagination Framework (Zhang et al., 2023b). We use the same predictive model architecture (ResNet50 trained from scratch), generate the same quantity of synthetic data (30×30$\times$30 × dataset size for Pets, 20×20$\times$20 × dataset size for Caltech101, Cars, and Flowers), the same diffusion model (Stable Diffusion v1.4), the same diffusion sampler (DDIM), and the same number of sampling iterates ($50$). We reproduce baseline models (not exposed to synthetic data or data augmentation conditions) to match the performance of GIF baselines (Original row in Table 1) to within $1\%$ of the generalization performance reported in GIF. We then fine-tune for $100$ epochs, generating synthetic data with Longtail Guidance according to the schedule in Table 5 (e.g. generate synthetic data in epoch 0, fine-tune until epoch 5, generate more synthetic data, fine-tune until epoch 10, …). We experimentally found that synthetic data iteratively generated throughout fine-tuning outperformed synthetic data generated all at once and speculate that this is due to the model-based longtail evolving throughout training. We also experimentally found that using total uncertainty (epistemic + aleatoric, the first term in the RHS of Equation 4) as the LTG guidance signal from the Epistemic Head was overall slightly more performant for downstream predictive model generalization improvements (as compared to epistemic or aleatoric alone), likely because several of our benchmark datasets involve fine-grained distinctions that benefit from additional synthetic data, not just towards the edge of the class manifold (epistemic) but also along decision boundaries (aleatoric). See Figure 4 for visualization.
@@ A.3 Dataset Details
A.4 Can We Continuously Mine for Additional High-Value Synthetic Data?
@@ A.6 Longtail Signals
A.7 Why Does Longtail Guidance Work without Training the Predictive Model on Intermediate Diffusion States?
A key finding of this work is that an existing predictive model $f_{\phi}$ does not need to be retrained on intermediate, noisy diffusion states to effectively guide diffusion model $\epsilon_{\theta}$ towards high-value synthetic training examples that are rare or hard from $f$’s perspective. This realization frees us from the dilemma of having to decide whether to waste predictive model capacity training on intermediate diffusion data it will never see in production or risking divergence from the production model by fine-tuning on intermediate noisy states. But why does this work?
In Figure  15, we visualize the intermediate, noisy diffusion states of two quantities: the decoded, predicted terminal state $D(P(\hat{x}_{t}^{0}))$ (top row, what LTG uses as guidance input to predictive model $f_{\phi}$) and the decoded data state, $D(x_{t})$ (bottom row, what classifier guidance would traditionally performs guidance on if not in a latent space). We observe that the terminal state predictions much more readily resemble natural image data at a much earlier time in the diffusion process (within the first 10% of denoising steps) than do the data states. In fact, decoded data states have off-distribution noise artifacts up through the first 90% of the diffusion denoising steps. We speculate that one reason the original classifier guidance work (Dhariwal & Nichol, 2021) performed guidance on intermediate noisy states is that they can be more efficiently generated than the predicted clean terminal states; intermediate states merely require sampling a random clean data point, sampling a random timestep, and applying the noise schedule. Predicted clean terminal states from a given denoising timestep additionally require a forward pass from the denoising network (see Equation 8).
@@
Results are plotted in Figure 15. Observe that the decoded, predicted terminal state $D(P(\hat{x}_{t}^{0}))$ are dramatically closer in distribution to real training data than are the naively decoded data states, $x_{t}$ for nearly all denoising iterations.
@@ A.8 Ablations
A.9 Additional Examples
In Figure 18, we show high-resolution examples of synthetic data generated by baseline diffusion and by Longtail Guidance. Guidance is performed by a SOTA ViT-based ImageNet-LT classifier described in Section 4.2.
Figure 18: Additional example of baseline diffusion (left) and Longtail Guidance (right). Baseline diffusion tends to generate canonical, well-posed views that help a predictive model generalize, but only up to a point. In contrast, Longtail Guidance produces more extreme, occluded, or challenging views from the perspective of a predictive model, enabling significantly improved generalization performance. In order, classes are: jellyfish, flamingo, lion, monarch butterfly, ant, candle, jeep, forklift, and ambulance.